a.) Find the point of intersection between the lines :<3, - 1,2> +<1, 1, - 1> and <-8, 2, 0> +t <-3,2-7>.

b.) show that the lines x +1 =3t, y=1, z +5 = 2t for t€ R and x +2 =s, y-3 = - 5s, z +4=-2s for t € R intersect, and find the point of intersection.

c.) Find the point of intersection between the planes : - 5x + y - 2z =3 and 2x - 3y +5z =-7.


D.)let L be the line given by <3, - 1,2> +t<1,1-1>, for t € R.

1.) show that the above line L lies on the plane - 2x + 3y - 4z +1 =0

2.)Find an equation for the plane through the point P =(3, - 2,4)that is perpendicular to the line <-8, 2, 0> +t<-3,2,-7>



1
Expert's answer
2021-07-05T16:45:40-0400

a)


"3+s=-8-3t""-1+s=2+2t""2-s=0-7t"


"s=-3t-11""4=-5t-10""1=-5t+2"



"s=-3t-11""3=-12""1=-5t+2"

No solution.

There is no point of intersection.


b)


"3t-1=s-2""1=-5s+3""2t-5=-2s-4"


"s=3t+1""5s=2""2t=-2s+1"



"t=-\dfrac{1}{5}""s=\dfrac{2}{5}""t=\dfrac{1}{10}"


No solution.

There is no point of intersection.


с)


"-5x+y-2z=3""2x-3y+5z=-7"



"-13x-z=2""2x-3y+5z=-7"

"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{2}{3}x+\dfrac{5}{3}z+\dfrac{7}{3}"



"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{29}{13}+\dfrac{21}{13}z"

"z\in\R"

The intersection between the planes is the line :<3, - 1,2> +<1, 1, - 1>


"\langle-\dfrac{2}{13}, \dfrac{29}{13}, 0\rangle+t\langle-\dfrac{1}{13}, \dfrac{1}{13},1\rangle, t\in \R"

d)

1)


"-2x+3y-4z+1"

"-6-2t-3+3t-8+4t+1"

"=5t-16\not=0\ for\ t=0"

The line does not lie on the plane.


2) Line


"\dfrac{x+8}{-3}=\dfrac{y-2}{2}=\dfrac{z-0}{-7}"

Point "P=(3, -2, 4)"

The equation for the plane through the point P that is perpendicular to the line is


"-3(x-3)+2(y-(-2))-7(z-4)=0"

"-3x+9+2y+4-7z+28=0"

"3x-2y+7z-41=0"